# Delay-Aware Robust Edge Network Hardening Under Decision-Dependent Uncertainty

### v1

@@ Abstract
Edge computing promises to offer low-latency and ubiquitous computation to numerous devices at the network edge. For delay-sensitive applications, link delays can have a direct impact on service quality. These delays can fluctuate drastically over time due to various factors such as network congestion, changing traffic conditions, cyberattacks, component failures, and natural disasters. Thus, it is crucial to efficiently harden the edge network to mitigate link delay variation as well as ensure a stable and improved user experience. To this end, we propose a novel robust model for optimal edge network hardening, considering the link delay uncertainty. Departing from the existing literature that treats uncertainties as exogenous, our model incorporates an endogenous uncertainty set to properly capture the impact of hardening and workload allocation decisions on link delays. However, the endogenous set introduces additional complexity to the problem due to the interdependence between decisions and uncertainties. We present two efficient methods to transform the problem into a solvable form. Extensive numerical results are shown to demonstrate the effectiveness of the proposed approach.
@@ I Introduction
Effectively managing uncertainties is a key factor for achieving consistent performance, reliability, and an enhanced user experience in EC. A crucial but frequently overlooked aspect in the literature is link delay uncertainty, which may arise due to a variety of factors, including fluctuating workloads, network congestion, routing changes, link failures, and cyber-attacks. Specifically, workload is a major contributor to delay uncertainty, as higher traffic volumes on a link can lead to increased packet delays. Network congestion is another critical factor that can cause delays and increased packet loss, particularly during high-volume network traffic periods. Moreover, routing changes, prompting packets to follow different paths, can introduce additional delays and variability in packet transmission. In the event of link failures, congestion can arise, causing slower speeds as traffic is rerouted. Finally, deliberate network attacks targeting links can inundate the network with excessive traffic, thereby exacerbating delay issues.
@@
Delay uncertainties can significantly affect the performance of edge network systems, especially for delay-sensitive applications. To address this issue, this paper focuses on enhancing the resilience of edge networks against delay uncertainties through various link hardening strategies, such as redundancy and failover, load balancing, traffic shaping, fault-tolerant routing, quality of service (QoS), and security mechanisms. Redundancy ensures that backup links and resources are available in the event of primary link failures [3, 4], while failover mechanisms automatically switch to backup links to guarantee network continuity [5]. Upgrading hardware, increasing bandwidth, and investing in higher-quality fiber optic cables can also bolster the resilience of network links. Load balancing is another crucial strategy that evenly distributes network traffic across multiple links, preventing any single link from becoming overloaded and reducing the likelihood of delays and failures [6]. Traffic shaping plays a vital role in controlling the traffic flow, avoiding link overload, and reducing delays and packet loss [7]. Fault-tolerant routing algorithms can help ensure that traffic continues to flow even when links fail, by rerouting traffic through alternate paths [2]. In addition, implementing QoS mechanisms and security measures is essential. QoS mechanisms prioritize traffic based on its importance, ensuring that critical traffic receives preferential treatment and minimizing delays and packet loss for critical applications. Security measures such as firewalls, intrusion detection systems, and other security mechanisms protect network links from attacks and unauthorized access, further fortifying the network [8, 9, 10].
@@
Different levels of hardening strategies yield varying degrees of network resilience [11]. Higher levels of hardening offer more robust protection, reducing the chance of network failures and improving overall network reliability [12]. However, these advantages are accompanied by increased costs. The associated costs differ due to each hardening level potentially requiring distinct technologies, strategies, infrastructure enhancements, or ongoing maintenance efforts. For example, standard redundancy and failover mechanisms provide essential protection against link failures, but they may fall short for critical, delay-sensitive services. Advanced hardening strategies (e.g., high-grade encryption, fault-tolerance routing, and extensive traffic shaping) provide a higher level of security and resilience [2]. Implementing these advanced measures often requires significant investments in software development and cybersecurity expertise [13, 14]. Also, establishing redundant high-capacity links, such as fiber optic cables, involves considerable capital outlay for infrastructure and continuous operational costs.
@@
This paper aims to bridge this gap by proposing a novel robust optimization (RO) framework for optimal link hardening in EC, considering the influence of the hardening and workload allocation decisions on link delays. Specifically, to accurately model the link delays, we introduce a decision-dependent uncertainty (DDU) set that explicitly captures the interdependence between the uncertainties and decisions. The incorporation of the DDU set into the robust model provides a more realistic representation of uncertainty during actual operation. Unlike the traditional RO method using a fixed decision-independent uncertainty (DIU) set, the DDU set is adjustable. This allows decision-makers to proactively control the form and level of uncertainty through their decisions, thereby mitigating the inherent conservatism of the robust solution. However, the introduction of interdependencies between uncertainties and decisions adds complexity to the problem, resulting in a large-scale non-linear optimization problem with numerous bilinear terms. To overcome this challenge, we present two efficient exact algorithms that can solve the underlying robust model under DDU. Our main goal is to quantify the benefit and importance of considering the endogenous uncertainty. Our contributions can be summarized as follows:
@@
  • Modeling: We introduce a new robust model to address the problem of optimal edge network link hardening and workload allocation, aiming to minimize the cost while enhancing user experience by reducing link delay variation. Different from traditional RO models under exogenous uncertainty, our proposed model incorporates an endogenous uncertainty set to capture the interdependencies between uncertain link delays and the decisions related to hardening and workload allocation. To the best of our knowledge, this is the first robust link-hardening model in EC, which explicitly accounts for this decision-dependent uncertainty.
@@
  • Techniques: To tackle the challenging robust problem under DDU, we depart from traditional RO frameworks and develop two efficient and exact reformulations, termed RDDU and e-RDDU. These reformulations are achieved through a series of transformations that convert the problem into a Mixed Integer Linear Programming (MILP) form, which can be solved efficiently using widely available solvers such as Gurobi [^1] https://www.gurobi.com/ and Mosek [^2] https://www.mosek.com/. Notably, the e-RDDU reformulation takes advantage of the specific structure of the DDU set and employs a clever transformation technique, which helps substantially reduce the size and complexity of the problem.
@@
  • Numerical results: Through extensive simulations, we demonstrate the efficiency of the proposed scheme compared to three benchmarks: a no-hardening scheme, a random hardening scheme, and a stochastic scheme under decision-dependent uncertainty. Additionally, we conduct sensitivity analyses to evaluate the impact of important system parameters on the system performance.
@@
The rest of the manuscript is organized as follows. Section  II introduces the system model and formulates the problem, while Section III expounds on the proposed solution approach. The numerical results are presented in Section IV. Section V discusses related work followed by conclusions in Section VI.
@@ II-A System Model
This paper studies an edge network hardening problem that involves a budget-constrained EC platform, aiming to optimize link hardening decisions to improve user experience by reducing delays. The platform manages a set $\mathcal{J}$ of $J$ heterogeneous ENs and offers edge resources to users in different areas, each represented by an access point (AP). The set of APs is denoted by $\mathcal{I}$, and the number of APs is $I$. Let $i$ and $j$ signify the AP index and EN index, respectively. We consider the graph $\mathbb{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes including $I$ APs and $J$ ENs. The set =$\mathcal{E}=\{(i,j):i\in\mathcal{I},j\in\mathcal{J}\}$ represents the set of links connecting APs with ENs. Each link $(i,j)$ represents the logical link connecting area $i$ with EN $j$, which may encompass multiple physical links.
@@
The quality of service for latency-sensitive applications is heavily affected by link delays. The delay in the logical link between area $i$ and EN $j$, which may encompass multiple physical link segments connecting intermediate nodes along the path between them, is denoted by $d_{i,j}$. Link delay can fluctuate dramatically over time due to various factors such as varying traffic conditions, cyberattacks, network congestion, and link/router failures. We consider $d_{i,j}$ as a time-varying and uncertain parameter instead of a constant as commonly assumed in the literature. To provide a superior user experience, the platform must aim to reduce not only the link delay but also the delay variation. Thus, developing efficient models for optimal link hardening that can mitigate delay variation and enhance system resilience is of paramount importance.
@@
The link hardening cost can vary based on several factors, such as the links’ characteristics and locations, as well as the types and levels of hardening. We assume that each link has $R$ levels of hardening, denoted by $h_{i,j}^{r}$ for level-$r$ hardening cost of link $(i,j)$, which increases as the hardening level increases. A higher hardening level implies improved link performance and reduced delay variation. We introduce a binary variable $t_{i,j}^{r}$, where $t_{i,j}^{r}$ equals $1$ if link $(i,j)$ if level-$r$ hardening is applied to link $(i,j)$. We also denote the incremental link hardening cost between two adjacent levels for each link by $\Delta h_{i,j}^{r}$ (i.e., $\Delta h_{i,j}^{r}=h^{r}_{i,j}-h^{r-1}_{i,j}$) and the incremental impact factor between adjacent levels by $\Delta\gamma_{i,j}^{r}$ (i.e., $\Delta\gamma_{i,j}^{r}=\gamma^{r}_{i,j}-\gamma^{r-1}_{i,j}$). Without loss of generality, we assume $\Delta h_{i,j}^{r}=\Delta h$ and $\Delta\gamma_{i,j}^{r}=\Delta\gamma,\forall i,j,r$. Additionally, let $B$ be the total hardening budget of the platform. The platform aims to identify a subset of critical links for efficient hardening within the budget constraint. The main notations are summarized in Table I.
II-B Deterministic Hardening Model
The EC platform aims to minimize the total link hardening cost while enhancing user experience by reducing the overall network delay and unmet demand. Without uncertainty consideration, the deterministic problem for optimizing the hardening and workload allocation decisions can be expressed as:
DET: $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}\>\mathcal{C}_{1}+\mathcal{C}_{2}$ (1a)
[…]
where $\mathcal{C}_{1}$ represents the total hardening cost and $\mathcal{C}_{2}$ represents the total penalty cost. The objective function and constraints are described in the following.
Objective function: The first term in the objective function (1a) represents the total link hardening cost given as follows:
$\mathcal{C}_{1}=\sum_{i,j}h_{i,j}^{r}t_{i,j}^{r}.$ (2)
The second term captures the delay penalty and unmet demand penalty from the workload allocation decision, i.e., we have:
$\mathcal{C}_{2}=\sum_{i,j}\rho d_{i,j}x_{i,j}+\sum_{i}s_{i}w_{i},$ (3)
where $\rho$ is the delay penalty parameter, which can be adjusted by the platform. A larger $\rho$ indicates the platform gives more priority to delay reduction than to unmet demand reduction.
Budget constraint: The investment budget of the platform is often limited. Consequently, the total link hardening cost cannot surpass this budget. This leads to the following constraint:
$\sum_{i,j,r}h_{i,j}^{r}t_{i,j}^{r}\leq B.$ (4)
Hardening constraints: Since the platform can choose at most one hardening level for each link $(i,j)$, we have:
$\sum_{r}t_{i,j}^{r}\leq 1,~{}\forall i,j.$ (5)
Capacity constraints: The total workload assigned to each EN cannot exceed its capacity, which can be expressed as follows:
$0\leq\sum_{i}x_{i,j}\leq C_{j},~{}\forall j.$ (6)
Workload allocation constraints: The demand from each area must be either served by some ENs or dropped, which can be captured by the following constraints:
$w_{i}+\sum_{j}x_{i,j}=\lambda_{i},~{}\forall i.$ (7)
QoS constraints: For quality control, the platform may impose a restriction that the proportion of unmet demand in any given area $i$ should not exceed a given threshold $\alpha_{i}$ as shown below:
$\frac{w_{i}}{\lambda_{i}}\leq\alpha_{i},~{}\forall i.$ (8)
Delay constraints: To ensure a satisfactory user experience, the platform can impose a certain threshold $\Delta_{i}$ on the average network delay for each area $i$. Note that $\frac{x_{i,j}}{\lambda_{i}}$ represents the proportion of workload from area $i$ allocated to EN $j$. The expression for this relationship is as follows:
$\sum_{j}d_{i,j}\frac{x_{i,j}}{\lambda_{i}}\leq\Delta_{i},~{}\forall i.$ (9)
Decision variables: Decision variables include the workload allocation ${\bm{x}}$, unmet demand ${\bm{w}}$ and hardening decision ${\bm{t}}$.
${\bm{w}}\in\mathbb{Z}_{+}^{I},{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},{\bm{t}}\in\{0,1\}^{I\times J\times R}.$ (10)
In practice, the platform typically provisions a discrete number of resource units (e.g., vCPUs, virtual machines) at each EN. Thus, the resource demands, allocation quantities, and unmet demands are represented as integers in our model.
@@ B Total link hardening budget
$t_{i,j}$$\{0,1\}$, 1 if link $(i,j)$ is chosen for level-$r$ hardening
@@
Remark. In the deterministic model, the link delay parameter $d_{i,j}$ is a constant. Also, the hardening decision has no impact on the link delay. Hence, from (1), it is easy to see that the optimal solution for ${\bm{t}}$ is $t_{i,j}=0,~{}\forall i,j$. Then, the deterministic model DET can be simplified as:
NH: $\min_{{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},{\bm{w}}\in\mathbb{Z}_{+}^{I}}\>\mathcal{C}_{2}$ (11)

### v2

@@ Abstract
Edge computing promises to offer low-latency and ubiquitous computation to numerous devices at the network edge. For delay-sensitive applications, link delays significantly affect service quality. These delays can fluctuate substantially over time due to various factors such as network congestion, changing traffic conditions, cyberattacks, component failures, and natural disasters. Thus, it is crucial to efficiently harden the edge network to mitigate link delay variation and ensure a stable and improved user experience. To this end, we propose a novel robust model for optimal edge network hardening, considering link delay uncertainty. Unlike existing literature that treats uncertainties as exogenous, our model incorporates an endogenous uncertainty set to properly capture the impact of hardening and workload allocation decisions on link delays. However, the endogenous set introduces additional complexity to the problem due to the interdependence between decisions and uncertainties. To address this, we present two efficient methods to transform the problem into a solvable form. Extensive numerical results demonstrate the effectiveness of the proposed approach in mitigating delay variations and enhancing system performance.
@@ I Introduction
Effectively managing uncertainties is a key factor for achieving consistent performance, reliability, and an enhanced user experience in EC. A crucial but frequently overlooked aspect in the literature is link delay uncertainty, which may arise due to a variety of factors, including fluctuating workloads, network congestion, routing changes, link failures, and cyber-attacks. Specifically, workload is a major contributor to delay uncertainty, as higher traffic volumes on a link can increase packet delays. Network congestion is another critical factor, causing delays and packet loss, particularly during high-traffic periods. Moreover, routing changes, prompting packets to follow different paths, can introduce additional delays and variability in packet transmission. In the event of link failures, congestion can arise, causing slower speeds as traffic is rerouted. Finally, deliberate network attacks targeting links can inundate the network with excessive traffic, thereby exacerbating delay issues.
@@
Delay uncertainties can significantly affect the performance of edge network systems, especially for delay-sensitive applications. To address this issue, this paper focuses on enhancing the resilience of edge networks against delay uncertainties through various link hardening strategies, such as redundancy and failover, load balancing, traffic shaping, fault-tolerant routing, quality of service (QoS), and security mechanisms. Redundancy ensures that backup links and resources are available in the event of primary link failures [3, 4], while failover mechanisms automatically switch to backup links to guarantee network continuity [5]. Upgrading hardware, increasing bandwidth, and investing in higher-quality fiber optic cables can also bolster the resilience of network links. Load balancing is another crucial strategy that evenly distributes network traffic across multiple links, preventing any single link from becoming overloaded and reducing the likelihood of delays and failures [6].
Traffic shaping plays a vital role in controlling the traffic flow, avoiding link overload, and reducing delays and packet loss [7]. Fault-tolerant routing algorithms can help ensure that traffic continues to flow even when links fail, by rerouting traffic through alternate paths [2]. In addition, implementing QoS mechanisms and security measures is essential. QoS mechanisms prioritize traffic based on its importance, ensuring that critical traffic receives preferential treatment and minimizing delays and packet loss for critical applications. Security measures such as firewalls, intrusion detection systems, and other security mechanisms protect network links from attacks and unauthorized access, further fortifying the network [8, 9, 10].
@@
Different levels of hardening strategies yield varying degrees of network resilience, each with its distinct cost [11]. Each level may contain one or several strategies, and the operator can flexibly choose which strategies to apply at each level for each link. Higher levels of hardening offer more robust protection, reducing the chance of network failures and improving overall network reliability [12]. However, these advantages are accompanied by increased costs. The associated costs differ due to each hardening level potentially requiring distinct technologies, strategies, infrastructure enhancements, or ongoing maintenance efforts. For example, standard redundancy and failover mechanisms provide essential protection against link failures, but they may fall short for critical, delay-sensitive services. Advanced hardening strategies (e.g., high-grade encryption, fault-tolerance routing, and extensive traffic shaping) provide a higher level of security and resilience [2]. Implementing these measures often requires significant investments in software development and cybersecurity expertise [13, 14]. Establishing redundant high-capacity links, like fiber optic cables, also requires considerable capital and ongoing operational costs.
@@
This paper aims to bridge this gap by proposing a novel robust optimization (RO) framework for optimal link hardening in EC, considering the influence of the hardening and workload allocation decisions on link delays. Specifically, to accurately model the link delays, we introduce a decision-dependent uncertainty (DDU) set that explicitly captures the interdependence between the uncertainties and decisions. The incorporation of the DDU set into the robust model provides a more realistic representation of uncertainty during actual operation. Unlike the traditional RO method using a fixed decision-independent uncertainty (DIU) set, the DDU set is adjustable. This allows decision-makers to proactively control the form and level of uncertainty through their decisions, thereby mitigating the inherent conservatism of the robust solution. However, the introduction of interdependencies between uncertainties and decisions adds complexity to the problem, resulting in a large-scale non-linear optimization problem with numerous bilinear terms. To overcome this challenge, we present two efficient exact algorithms that can solve the underlying robust model under DDU. Our main goal is to quantify the benefit and importance of considering endogenous uncertainty. The proposed model is primarily designed for edge infrastructure providers that own and operate distributed edge data centers and network links. For example, telecom companies such as Verizon, T-Mobile, and AT&T can benefit directly from its implementation. Additionally, cloud and edge infrastructure providers (e.g., Amazon, Google, and Equinix), which manage extensive network infrastructures, can also leverage the model.
Our contributions can be summarized as follows:
@@
  • 1) Modeling: We introduce a novel robust model that addresses the optimal edge network link hardening and workload allocation problem, aiming to minimize costs while enhancing user experience by reducing link delay variation. Unlike traditional RO models, which typically consider exogenous uncertainty [15], our proposed model incorporates an endogenous uncertainty set to capture the interdependencies between uncertain link delays and the decisions related to hardening and workload allocation. To the best of our knowledge, this is the first robust link-hardening model in EC that explicitly accounts for such decision-dependent uncertainty.
@@
  • 2) Techniques: To solve the challenging robust problem under DDU, we deviate from traditional RO models and develop two efficient and exact reformulations, termed RDDU and e-RDDU. These reformulations are derived through a series of transformations that convert the original problem into a Mixed Integer Linear Programming (MILP) form, which can be solved efficiently using standard solvers such as Gurobi [^1] https://www.gurobi.com/ and Mosek [^2] https://www.mosek.com/. Notably, e-RDDU enhances upon RDDU by leveraging the specific structure of the DDU set and employing a clever transformation technique that substantially reduces the problem’s size and complexity.
@@
  • 3) Numerical results: Extensive simulations demonstrate the efficiency of the proposed scheme compared to three benchmarks: a no-hardening scheme, a random hardening scheme, and a stochastic scheme under DDU. Also, sensitivity analyses were conducted to evaluate the impact of important system parameters on system performance.
@@
This paper is structured as follows. Section  II outlines the system model and problem formulation. Section III details the proposed solution approach. The numerical results are presented in Section IV. Section V discusses related work followed by conclusions in Section VI.
@@ II System Model and Problem Formulation
This section outlines the system model. We also define key parameters, decision variables, and constraints, and formulate the deterministic model, followed by two robust models that consider both exogenous and endogenous uncertainties.
@@ II-A System Model
This paper studies an edge network hardening problem that involves a budget-constrained EC platform (i.e., an edge infrastructure provider), aiming to optimize link hardening decisions to improve user experience by reducing delays. The platform manages a set $\mathcal{J}$ of $J$ heterogeneous edge nodes (ENs) and offers edge resources to users in different areas, each represented by an access point (AP). The set of APs is denoted by $\mathcal{I}$, and the number of APs is $I$. Let $i$ and $j$ signify the AP index and EN index, respectively. We consider the graph $\mathbb{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes including $I$ APs and $J$ ENs. The set =$\mathcal{E}=\{(i,j):i\in\mathcal{I},j\in\mathcal{J}\}$ represents the set of links connecting APs with ENs. Each link $(i,j)$ represents the logical link connecting area $i$ with EN $j$, which may encompass multiple physical links.
@@
The QoS for latency-sensitive applications is heavily affected by link delays. The delay in the logical link between area $i$ and EN $j$, which may encompass multiple physical link segments connecting intermediate nodes along the path between them, is denoted by $d_{i,j}$. Link delay can fluctuate dramatically over time due to various factors such as varying traffic conditions, cyberattacks, network congestion, and link/router failures. We consider $d_{i,j}$ as a time-varying and uncertain parameter instead of a constant as commonly assumed in the literature. To provide a superior user experience, the platform must aim to reduce both link delay and delay variation. Developing efficient models for optimal link hardening to mitigate delay variation and improve system resilience is of paramount importance.
@@
The cost of link hardening can vary depending on several factors, such as the characteristics and locations of the links, as well as the types and levels of hardening. We assume that each link has $R$ levels of hardening, denoted by $h_{i,j}^{r}$ for level-$r$ hardening cost of link $(i,j)$, which increases as the hardening level increases. A higher hardening level implies improved link performance and reduced delay variation. Note that each hardening level may contain several link hardening strategies. We introduce a binary variable $t_{i,j}^{r}$, where $t_{i,j}^{r}$ equals $1$ if link $(i,j)$ if level-$r$ hardening is applied to link $(i,j)$. We also denote the incremental link hardening cost between two adjacent levels for each link by $\Delta h_{i,j}^{r}$ (i.e., $\Delta h_{i,j}^{r}=h^{r}_{i,j}-h^{r-1}_{i,j}$) and the incremental impact factor (IF) between adjacent levels by $\Delta\gamma_{i,j}^{r}$ (i.e., $\Delta\gamma_{i,j}^{r}=\gamma^{r}_{i,j}-\gamma^{r-1}_{i,j}$). Without loss of generality, we assume $\Delta h_{i,j}^{r}=\Delta h$ and $\Delta\gamma_{i,j}^{r}=\Delta\gamma,\forall i,j,r$. Let $B$ be the total hardening budget of the platform. The platform aims to identify a subset of critical links for efficient hardening within budget constraints. Table I summarizes the main notations.
@@ B Total link hardening budget
$t_{i,j}^{r}$$\{0,1\}$, 1 if link $(i,j)$ is chosen for level-$r$ hardening
@@
Fig. 1 presents a toy example of the system model with $2$ AP and $2$ ENs. The platform aims to serve users’ requests from both areas. For simplicity, link conditions are represented as tuples and visualized using different colors. In scenario (a), the delay for link $(1,1)$ ranges from $[5,15]$ ms under normal conditions, while link $(1,2)$ experiences higher variability, ranging from $[5,20]$ ms. Without hardening, $\frac{3}{4}$ of the workload from AP1111 is allocated to EN1111, while $\frac{1}{4}$ is assigned to EN2222. The platform may priotize allocating workload to EN1111 if capacity is allowed due to its lower network delay. In scenario (b), proactive hardening is applied to link $(1,2)$, reducing the maximum delay deviation from $20$ ms to $12$ ms, thereby improving QoS. Consequently, the platform can priortize workload allocation to EN2222 due to its lower network latency. This example highlights the importance of proactive hardening in mitigating uncertainty, enhancing system performance, and reducing operational costs.
Figure 1: System model
II-B Deterministic Hardening Model
We first formulate the deterministic problem, termed DET, for optimizing hardening and workload allocation decisions without considering system uncertainties. The EC platform aims to minimize the total cost of link hardening while improving user experience by reducing network delay and unmet demand. The DET problem can be expressed as follows:
(DET) $\min_{{\bm{t}},{\bm{x}},{\bm{w}}}\>\underbrace{\sum_{i,j}h_{i,j}^{r}t_{i,j}^{r}}_{\mathcal{C}_{1}({\bm{t}})}+\underbrace{\sum_{i,j}\rho d_{i,j}x_{i,j}+\sum_{i}s_{i}w_{i}}_{\mathcal{C}_{2}({\bm{x}},{\bm{w}})}$ (1a)
s.t. $\sum_{i,j,r}h_{i,j}^{r}t_{i,j}^{r}\leq B,$ (1b)
$\sum_{r}t_{i,j}^{r}\leq 1,~{}\forall i,j,$ (1c)
$0\leq\sum_{i}x_{i,j}\leq C_{j},~{}\forall j.$ (1d)
$w_{i}+\sum_{j}x_{i,j}=\lambda_{i},~{}\forall i,$ (1e)
$\frac{w_{i}}{\lambda_{i}}\leq\alpha_{i},~{}\forall i,$ (1f)
$\sum_{j}d_{i,j}\frac{x_{i,j}}{\lambda_{i}}\leq\Delta_{i},~{}\forall i,$ (1g)
${\bm{w}}\in\mathbb{Z}_{+}^{I},~{}{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},~{}{\bm{t}}\in\{0,1\}^{I\times J\times R}.$ (1h)
where $\mathcal{C}_{1}({\bm{t}})$ represents the total hardening cost and $\mathcal{C}_{2}({\bm{x}},{\bm{w}})$ represents the total penalty cost, including both the delay penalty and the unmet demand penalty resulting from workload allocation decisions ${\bm{x}}$ and ${\bm{w}}$. Note that $\rho$ is the delay penalty parameter controlled by the platform. A higher value of $\rho$ indicates that the platform prioritizes reducing delays over reducing unmet demand. The underlying optimization problem of the platform is subject to constraints (1b)-(1h).
Given that the platform’s investment budget is often limited, constraint (1b) ensures that the total link hardening cost cannot exceed the budget. The platform can choose at most one hardening level for each link $(i,j)$, as shown in (1c) (see Remark II.2). Constraints (1d) specify that the total workload assigned to each EN cannot exceed its capacity. The demand from each area $i$ must be served by some EN (${\bm{x}}$) or dropped (${\bm{w}}$), as captured by (1e). For quality control, constraints (1f) enforce that the proportion of unmet demand in any area $i$ does not exceed a predefined threshold $\alpha_{i}$. Additionally, to maintain a satisfactory user experience, constraints (1g) impose a threshold $\Delta_{i}$ on the average network delay for each area $i$. Here, $\frac{x_{i,j}}{\lambda_{i}}$ represents the proportion of workload from area $i$ allocated to EN $j$. Finally, constraints (1h) define all decision variables to be optimized in the model, where resource demands, allocation quantities, and unmet demands are represented as integers, as the platform typically provisions a discrete number of resource units (e.g., vCPUs, virtual machines) at each EN.
Remark II.1.
In the deterministic model, the link delay parameter $d_{i,j}$ is a constant. Also, the hardening decision has no impact on the link delay. Hence, from (1), it is easy to see that the optimal solution for ${\bm{t}}$ is $t_{i,j}=0,~{}\forall i,j$. Then, the deterministic model DET can be simplified as:
(NH) $\min_{{\bm{x}}\in\mathbb{Z}_{+}^{I\times J},{\bm{w}}\in\mathbb{Z}_{+}^{I}}\>\mathcal{C}_{2}$ (2)